OpenStudy (anonymous):

Integral (2-3x) cos9xdx

OpenStudy (mathteacher1729):

You can break this up like this: \[\int2\cos(9x)dx - 3\int x\cos(9x)dx\] The first integral should be easy to compute, the second will involve integration by parts. http://tutorial.math.lamar.edu/Classes/CalcII/IntegrationByParts.aspx Hope this helps.

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

\[\int\limits_{ }^{}\cos9xdx=1/9 \sin9x\] \[\int\limits_{ }^{}x*\cos 9x dx=\] u=x du=dx dv=cos 9x v=1/9 sin9x (integration by parts) then: \[\int\limits_{ }^{}x* \cos9x = 1/9 *x*\sin9x - 1/9 \int\limits_{ }^{}\sin9x dx=\] can you finish it?

OpenStudy (anonymous):

yea thanks, i have done it using integration by part, u = (2-3x) dv = cos9xdx

OpenStudy (anonymous):

awesome!